OpenStudy (anonymous):

Find the slope and the y-intercept of the equation x − 4(y + 1) − 5 = 0.

OpenStudy (amistre64):

let x=0, solve for y .... that is the intercept let x=1, solve for y again ... the difference the y values from x=0 to x=1 will define the slope.

OpenStudy (anonymous):

Hi Look : if we have : 2*(2+3)=2*2+2*3 or : 2(a+8)=2*a+2*8=2a+16 k ? So : − 4(y + 1) = (-4)*y+(-4)*+1=-4y-4 ( look : (-)*(-)=+ (+)*(+)=+ (_)*(+)=- ) so we have : x − 4(y + 1) − 5 = 0=>x-4y-4-5=0 Get it in standard form . it s : y=mx+d ;so first we should have y alone : x-4-5=4y x-9=4y delete 4 in 4y ; divided 4 x/4-9/4=y=>1/4x-9/4=y m x + d = y So we have : slope : that number you see before x means 1/4 y-intercept :is that -9/4 Got it ?;) @dfghjkutre